# Multidimensionality of the Hubble tension: the roles of $\Omega_{m}$ and $\omega_{c}$

### v1

@@ Abstract
The Hubble tension is inherently multidimensional, and bears important implications for parameters beyond $H_{0}$. We discuss the key role of the matter density parameter $\Omega_{m}$ and the physical cold dark matter density $\omega_{c}$. We argue that once $\Omega_{m}$ and the physical baryon density $\omega_{b}$ are calibrated, through Baryon Acoustic Oscillations (BAO) and/or Type Ia Supernovae (SNeIa) for $\Omega_{m}$, and via Big Bang Nucleosynthesis for $\omega_{b}$, any model raising $H_{0}$ inevitably raises $\omega_{c}$ and, under minimal assumptions, the clustering parameter $S_{8}$. We explicitly verify that this behaviour holds when analyzing recent BAO and SNeIa data. We argue that a calibration of $\Omega_{m}$ which is as reliable and model-independent as possible should be a priority in the Hubble tension discussion, and an interesting possibility in this sense could be represented by galaxy cluster gas mass fraction measurements.
@@ I Introduction
The ability of the simple, six-parameter $\Lambda$CDM cosmological model to account for a wide range of cosmological observations remains the main reason for its continued success. Yet, with the increase in precision of cosmological observations over the past decade, this success has been challenged by various tensions whose significance keeps growing: an example is the mismatch between the value of the Hubble constant $H_{0}=(67.36\pm 0.54)\,{\text{km}}/{\text{s}}/{\text{Mpc}}$ inferred from Cosmic Microwave Background (CMB) data within $\Lambda$CDM Aghanim et al. (2020), and a number of direct measurements of the same quantity, including but not limited to the SH0ES Cepheid-calibrated Type Ia Supernovae (SNeIa) local distance ladder measurement of $H_{0}=(73.04\pm 1.04)\,{\text{km}}/{\text{s}}/{\text{Mpc}}$ Riess et al. (2022), see e.g. Refs. Verde et al. (2019); Di Valentino et al. (2021a, b); Perivolaropoulos and Skara (2022a); Schöneberg et al. (2022a); Shah et al. (2021); Abdalla et al. (2022); Di Valentino (2022); Hu and Wang (2023); Verde et al. (2023) for reviews. Explanations based on systematics (e.g. Refs. Mortsell et al. (2022a, b); Wojtak and Hjorth (2022)) are growing ever more challenging to defend, which is why serious consideration has been given to the possibility that the Hubble tension calls for new physics beyond $\Lambda$CDM (with no claims as to completeness, see Refs. Anchordoqui et al. (2015); Karwal and Kamionkowski (2016); Zhao et al. (2017); Benetti et al. (2018); Mörtsell and Dhawan (2018); Vagnozzi et al. (2018); Kumar et al. (2018); Yang et al. (2018); Banihashemi et al. (2020); Guo et al. (2019); Graef et al. (2019); Banihashemi et al. (2019); Agrawal et al. (2023); Li and Shafieloo (2019); Yang et al. (2019); Vagnozzi (2020); Visinelli et al. (2019); Di Valentino et al. (2020a); Escudero and Witte (2020); Di Valentino et al. (2020b); Niedermann and Sloth (2021a); Sakstein and Trodden (2020); Ye and Piao (2020); Hogg et al. (2020); Ballesteros et al. (2020); Alestas et al. (2020); Jedamzik and Pogosian (2020); Ballardini et al. (2020); Di Valentino et al. (2020c); Banerjee et al. (2021); Niedermann and Sloth (2020); Gonzalez et al. (2020); Braglia et al. (2021); Roy Choudhury et al. (2021); Brinckmann et al. (2021); Alestas et al. (2021); Gao et al. (2021); Alestas and Perivolaropoulos (2021); Karwal et al. (2022); Cyr-Racine et al. (2022); Akarsu et al. (2021); Niedermann and Sloth (2022); Saridakis et al. (2023); Sen et al. (2022); Herold et al. (2022); Odintsov and Oikonomou (2022a); Heisenberg et al. (2023, 2022); Sharma et al. (2022); Ren et al. (2022); Nunes et al. (2022); Nojiri et al. (2022); Schöneberg and Franco Abellán (2022); Joseph et al. (2023); Gómez-Valent et al. (2022); Moshafi et al. (2022); Odintsov and Oikonomou (2022b); Banerjee et al. (2023); Alvarez et al. (2024); Ge et al. (2023); Akarsu et al. (2023a); Gangopadhyay et al. (2023a); Schiavone et al. (2023); Gao et al. (2024); Brinckmann et al. (2023); Khodadi and Schreck (2023); Dahmani et al. (2023); Ben-Dayan and Kumar (2023); de Cruz Perez and Sola Peracaula (2024); Ballardini et al. (2023); Yao et al. (2024); Gangopadhyay et al. (2023b); Zhai et al. (2023); Sola Peracaula et al. (2023); Gómez-Valent et al. (2024); Ruchika et al. (2024a); Adil et al. (2024); Frion et al. (2023); Akarsu et al. (2023b); Escamilla et al. (2024); Petronikolou and Saridakis (2023); Sharma et al. (2024); Ben-Dayan and Kumar (2024); Ramadan et al. (2024); Fu and Wang (2024); Efstathiou et al. (2024); Montani et al. (2024a); Lazkoz et al. (2024); Forconi et al. (2024); Sebastianutti et al. (2024); Benisty et al. (2024); Stahl et al. (2024); Shah et al. (2024); Giarè et al. (2024); Montani et al. (2024b); Co et al. (2024); Akarsu et al. (2024a); Yadav et al. (2024); Dwivedi and Högås (2024); Montani et al. (2024c) for examples).
@@
Nevertheless, focusing on $H_{0}$ can only reveal part of the story and is, at best, misleading. A key role in the Hubble tension is in fact played by Baryon Acoustic Oscillation (BAO) measurements: once the sound horizon at baryon drag $r_{d}$ is calibrated to the $\Lambda$CDM value – either from the CMB or through the Big Bang Nucleosynthesis (BBN) determination of the physical baryon density $\omega_{b}$ – BAO can be combined with uncalibrated SNeIa measurements to build an inverse distance ladder. This yields a value of $H_{0}$ compatible with the “low” Planck $\Lambda$CDM one, yet in principle independent of any CMB data. Another way of seeing this is to note that the distance-redshift diagrams of SH0ES-calibrated SNeIa, and BAO calibrated through the $\Lambda$CDM value of $r_{d}$, are mutually completely inconsistent despite probing the same $0.1\lesssim z\lesssim 2$ redshift range (see e.g. Fig. 1 of Ref. Poulin et al. (2024), as well as Ref. Bousis and Perivolaropoulos (2024)). Given that BAO are sensitive to the combination $H_{0}r_{d}$, it becomes clear that solving the Hubble tension while not altering the SNeIa absolute magnitude $M_{B}$ necessarily requires a decrease in $r_{d}$. This, in turn, calls for new physics operating before recombination Bernal et al. (2016); Addison et al. (2018); Lemos et al. (2019); Aylor et al. (2019); Schöneberg et al. (2019); Knox and Millea (2020); Arendse et al. (2020); Efstathiou (2021); Cai et al. (2022); Keeley and Shafieloo (2023) [^1] See Ref. Akarsu et al. (2023c) for possible caveats regarding this conclusion.
@@
However, even looking at $H_{0}$ and $r_{d}$ is still not the end of the story, as models that aim to solve the tension are inevitably subject to several other constraints. Lowering $r_{d}$, therefore, is just a part of what a successful model should do, and various other cosmological quantities such as the age of the Universe $t_{U}$ Jimenez et al. (2019); Bernal et al. (2021); Vagnozzi et al. (2022); Wei and Melia (2022); Capozziello et al. (2023), the physical matter density $\omega_{m}$ Pogosian et al. (2020); Jedamzik et al. (2021); Bernal et al. (2021), and the fractional matter density parameter $\Omega_{m}$ Lin et al. (2020, 2021); Bernal et al. (2021); Sakr (2023) appear to play a key role in the Hubble tension discourse. It has therefore become increasingly clear that the Hubble tension is inherently a multi-dimensional problem, as recently emphasized by Ref. Poulin et al. (2024), where the denomination “cosmic calibration tension” was suggested.
@@
Our work fits in this context by further exploring the role played by two important cosmological parameters: the fractional matter density parameter $\Omega_{m}$, and the physical dark matter (DM) density parameter $\omega_{c}=\Omega_{c}h^{2}$, with $h$ the dimensionless Hubble constant. Assuming both $\Omega_{m}$ and the physical baryon density $\omega_{b}$ are somehow calibrated, one can easily show that an increase in $H_{0}$ must be accompanied by an increase in $\omega_{c}$. Under minimal assumptions concerning the primordial power spectrum, this inevitably leads to a potentially problematic increase in the clustering parameter $S_{8}$ Poulin et al. (2024). It so happens that $\Omega_{m}$ and $\omega_{b}$ can in fact be calibrated in a fairly robust way, through a combination of uncalibrated BAO and SNeIa in the former case, and BBN in the latter. In the rest of our work we discuss these points in more detail, explicitly demonstrating on real data how an increase in $H_{0}$ is accompanied by an increase in $\omega_{c}$ (and $S_{8}$), while studying the impact of the $\Omega_{m}$ calibration and arguing that a high-fidelity, as model-independent as possible calibration of $\Omega_{m}$, should be a major priority in the quest towards solving the Hubble tension.
@@ II Multidimensionality of the Hubble tension
$\frac{\delta\omega_{c}}{\omega_{c}}=2\frac{\delta\omega_{c}}{\omega_{m}}\frac{\omega_{m}}{\omega_{c}}\approx 2\left(1+\frac{\omega_{b}+\omega_{\nu}}{\omega_{c}}\right)\frac{\delta h}{h}\,.$ (4)
@@
We remark once more that the earlier arguments hinge upon calibrations for $\Omega_{m}$ and $\omega_{b}$. We now discuss in more detail how this is achieved. To set the stage, in what follows we work within a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric, and assume that the Etherington distance-duality relation (DDR) holds, in agreement with recent data Renzi et al. (2022). Within our assumptions, the transverse comoving distance $D_{M}$ is given by the following:
@@
A single BAO measurement (at a single effective redshift) of one among $\theta_{d}$, $\delta z_{d}$, or $\theta_{v}$ is unable to disentangle the effects of $r_{d}H_{0}$ and $\Omega_{m}$, which are completely degenerate between each other. However, this degeneracy can be partially broken if the BAO angular scale is measured over a sufficiently wide range of effective redshifts (at present, $0.1\lesssim z_{\text{eff}}\lesssim 2.5$). The reason is that the slope describing the $r_{d}H_{0}$-$\Omega_{m}$ correlation slowly changes with $z_{\text{eff}}$ (see e.g. Fig. 2 of Ref. Lin et al. (2021)), reflecting how the importance of the dark energy contribution relative to the matter one decreases with increasing $z_{\text{eff}}$. This is not sufficient to lead to a very precise inference of $\Omega_{m}$ at the current level of precision (which is why earlier we stated that the degeneracy is only partially broken), but illustrates where the sensitivity of BAO measurements to $\Omega_{m}$ arises from.
@@
For what concerns the physical baryon density $\omega_{b}$, this is tightly constrained by BBN considerations on the abundance of light elements, especially deuterium. This is because $\omega_{b}$ controls (among others) the neutron-to-proton density at the time of BBN, a parameter to which the yield of light elements is particularly sensitive. It is worth noting that the CMB is also sensitive to $\omega_{b}$ through its impact on the relative height of odd and even acoustic peaks. The signature of $\omega_{b}$ in CMB data is therefore quite clean, making its determination relatively stable across different models, while still being in part model-dependent. Nevertheless, in support of the BBN determination of $\omega_{b}$, it is worth noting that this is in excellent agreement with the $\Lambda$CDM-based CMB determination of the same parameter. Moreover, it has been shown in Ref. Motloch (2020) that $\omega_{b}$ can be inferred in eight independent ways from CMB data, reflecting eight independent ways in which the physical baryon density affects the CMB power spectra and with all eight determinations in agreement between each other. Therefore, while the earlier calibration of $\Omega_{m}$ is somewhat model-dependent (it implicitly depends on the assumed late-time background expansion), the calibration of $\omega_{b}$ can be considered extremely robust and model-independent.
@@
where $T(k)$ denotes the transfer function, $k_{\text{eq}}$ is the equality wavenumber, and $a_{\text{eq}}$ is the equality scale factor. This necessarily increases $S_{8}\sim H_{0}^{2}$ as well, worsening the discrepancy between values of $S_{8}$ determined by weak lensing Troxel et al. (2018); Heymans et al. (2021), redshift space distortion Efstathiou and Lemos (2018); Benisty (2021); Nunes and Vagnozzi (2021), and cluster counts Sakr et al. (2022), all of which fall somewhat lower compared to the Planck (TTTEEE+lowE+lensing) determination of $S_{8}=0.832\pm 0.013$ Aghanim et al. (2020) [^3] We note that recent analyses of cosmic shear data from DES Y3 and KiDS-1000 Abbott et al. (2023), and of the cluster mass function from SRG/eROSITA Ghirardini et al. (2024), have decreased the significance of the $S_{8}$ discrepancy. Analogously to Eq. (3), we can expect the following:

### v2

@@ Abstract
The Hubble tension is inherently multidimensional, and bears important implications for parameters beyond $H_{0}$. We discuss the key role of the matter density parameter $\Omega_{m}$ and the physical cold dark matter density $\omega_{c}$. We argue that once $\Omega_{m}$ and the physical baryon density $\omega_{b}$ are calibrated, through Baryon Acoustic Oscillations (BAO) and/or Type Ia Supernovae (SNeIa) for $\Omega_{m}$, and via Big Bang Nucleosynthesis for $\omega_{b}$, any model raising $H_{0}$ requires raising $\omega_{c}$ and, under minimal assumptions, also the clustering parameter $S_{8}$. We explicitly verify that this behaviour holds when analyzing recent BAO and SNeIa data. We argue that a calibration of $\Omega_{m}$ as reliable and model-independent as possible should be a priority in the Hubble tension discussion, and an interesting possibility in this sense could be represented by galaxy cluster gas mass fraction measurements.
@@ I Introduction
The ability of the simple, six-parameter $\Lambda$CDM cosmological model to account for a wide range of cosmological observations remains the main reason for its continued success. Yet, with the increase in precision of cosmological observations over the past decade, this success has been challenged by various tensions whose significance keeps growing: an example is the mismatch between the value of the Hubble constant $H_{0}=(67.36\pm 0.54)\,{\text{km}}/{\text{s}}/{\text{Mpc}}$ inferred from Cosmic Microwave Background (CMB) data within $\Lambda$CDM Aghanim et al. (2020), and a number of direct measurements of the same quantity, including but not limited to the SH0ES Cepheid-calibrated Type Ia Supernovae (SNeIa) local distance ladder measurement of $H_{0}=(73.04\pm 1.04)\,{\text{km}}/{\text{s}}/{\text{Mpc}}$ Riess et al. (2022), see e.g. Refs. Verde et al. (2019); Di Valentino et al. (2021a, b); Perivolaropoulos and Skara (2022a); Schöneberg et al. (2022a); Shah et al. (2021); Abdalla et al. (2022); Di Valentino (2022); Hu and Wang (2023); Verde et al. (2023) for reviews. Explanations based on systematics (e.g. Refs. Efstathiou (2020); Mortsell et al. (2022a, b); Wojtak and Hjorth (2022); Freedman et al. (2024)) are growing ever more challenging to defend, which is why serious consideration has been given to the possibility that the Hubble tension calls for new physics beyond $\Lambda$CDM (with no claims as to completeness, see Refs. Anchordoqui et al. (2015); Karwal and Kamionkowski (2016); Zhao et al. (2017); Benetti et al. (2018); Mörtsell and Dhawan (2018); Vagnozzi et al. (2018); Kumar et al. (2018); Yang et al. (2018); Banihashemi et al. (2020); Guo et al. (2019); Graef et al. (2019); Banihashemi et al. (2019); Agrawal et al. (2023); Li and Shafieloo (2019); Yang et al. (2019); Vagnozzi (2020); Visinelli et al. (2019); Di Valentino et al. (2020a); Escudero and Witte (2020); Di Valentino et al. (2020b); Niedermann and Sloth (2021a); Sakstein and Trodden (2020); Ye and Piao (2020); Hogg et al. (2020a); Ballesteros et al. (2020); Alestas et al. (2020); Jedamzik and Pogosian (2020); Ballardini et al. (2020); Di Valentino et al. (2020c); Banerjee et al. (2021); Niedermann and Sloth (2020); Gonzalez et al. (2020); Braglia et al. (2021); Roy Choudhury et al. (2021); Brinckmann et al. (2021); Alestas et al. (2021); Gao et al. (2021); Alestas and Perivolaropoulos (2021); Karwal et al. (2022); Cyr-Racine et al. (2022); Akarsu et al. (2021); Niedermann and Sloth (2022); Saridakis et al. (2023); Sen et al. (2022); Herold et al. (2022); Odintsov and Oikonomou (2022a); Heisenberg et al. (2023, 2022); Sharma et al. (2022); Ren et al. (2022); Nunes et al. (2022); Nojiri et al. (2022); Schöneberg and Franco Abellán (2022); Joseph et al. (2023); Gómez-Valent et al. (2022); Moshafi et al. (2022); Odintsov and Oikonomou (2022b); Banerjee et al. (2023); Alvarez et al. (2024); Ge et al. (2023); Akarsu et al. (2023a); Gangopadhyay et al. (2023a); Schiavone et al. (2023); Gao et al. (2024); Brinckmann et al. (2023); Khodadi and Schreck (2023); Dahmani et al. (2023); Ben-Dayan and Kumar (2023); de Cruz Perez and Sola Peracaula (2024); Ballardini et al. (2023); Yao et al. (2024a); Gangopadhyay et al. (2023b); Zhai et al. (2023); Sola Peracaula et al. (2023); Gómez-Valent et al. (2024); Ruchika et al. (2024a); Adil et al. (2024); Frion et al. (2023); Akarsu et al. (2023b); Escamilla et al. (2024a); Petronikolou and Saridakis (2023); Sharma et al. (2024); Ben-Dayan and Kumar (2024); Ramadan et al. (2024); Fu and Wang (2024); Efstathiou et al. (2024); Montani et al. (2024a); Lazkoz et al. (2024); Forconi et al. (2024); Sebastianutti et al. (2024); Benisty et al. (2024); Stahl et al. (2024); Shah et al. (2024); Giarè et al. (2024); Montani et al. (2024b); Co et al. (2024); Akarsu et al. (2024a); Yadav et al. (2024); Nozari et al. (2024); Dwivedi and Högås (2024); Montani et al. (2024c); Escamilla et al. (2024b); Perivolaropoulos (2024); Yao et al. (2024b) for examples).
@@
Nevertheless, focusing on $H_{0}$ can only reveal part of the story and is, at best, misleading. A key role in the Hubble tension is in fact played by Baryon Acoustic Oscillation (BAO) measurements: once the sound horizon at baryon drag $r_{d}$ is calibrated to the $\Lambda$CDM value – either from the CMB or through the Big Bang Nucleosynthesis (BBN) determination of the physical baryon density $\omega_{b}$ – BAO can be combined with uncalibrated SNeIa measurements to build an inverse distance ladder. This yields a value of $H_{0}$ compatible with the “low” Planck $\Lambda$CDM one, yet in principle independent of any CMB data. Another way of seeing this is to note that the distance-redshift diagrams of SH0ES-calibrated SNeIa, and BAO calibrated through the $\Lambda$CDM value of $r_{d}$, are mutually completely inconsistent despite probing the same $0.1\lesssim z\lesssim 2$ redshift range (see e.g. Fig. 1 of Ref. Poulin et al. (2024), Fig. 1 of Ref. Bousis and Perivolaropoulos (2024), as well as Refs. Raveri (2020); Benevento et al. (2020); Camarena and Marra (2021, 2023); Raveri (2023)). Given that BAO are sensitive to the combination $H_{0}r_{d}$, it becomes clear that solving the Hubble tension while not altering the SNeIa absolute magnitude $M_{B}$ necessarily requires a decrease in $r_{d}$. This, in turn, calls for new physics operating before recombination Bernal et al. (2016); Addison et al. (2018); Lemos et al. (2019); Aylor et al. (2019); Schöneberg et al. (2019); Knox and Millea (2020); Arendse et al. (2020); Efstathiou (2021); Cai et al. (2022); Keeley and Shafieloo (2023) [^1] See Refs. Akarsu et al. (2023c); Tutusaus et al. (2023); Escamilla et al. (2024c) for possible caveats regarding this conclusion.
@@
However, even looking at $H_{0}$ and $r_{d}$ is still not the end of the story, as models that aim to solve the Hubble tension are inevitably subject to several other constraints. Lowering $r_{d}$, therefore, is just a part of what a successful model should do, and various other cosmological quantities such as the age of the Universe $t_{U}$ Jimenez et al. (2019); Bernal et al. (2021); Vagnozzi et al. (2022); Wei and Melia (2022); Capozziello et al. (2023), the physical matter density $\omega_{m}$ Pogosian et al. (2020); Jedamzik et al. (2021); Bernal et al. (2021), and the fractional matter density parameter $\Omega_{m}$ Lin et al. (2020, 2021); Bernal et al. (2021); Sakr (2023); Baryakhtar et al. (2024) appear to play a key role in the Hubble tension discourse. It has therefore become increasingly clear that the Hubble tension is inherently a multi-dimensional problem, as recently emphasized by Ref. Poulin et al. (2024), where the denomination “cosmic calibration tension” was suggested.
@@
Our work fits in this context by further exploring the role played by two important cosmological parameters: the fractional matter density parameter $\Omega_{m}$, and the physical dark matter (DM) density parameter $\omega_{c}=\Omega_{c}h^{2}$, with $h$ the dimensionless Hubble constant. Assuming both $\Omega_{m}$ and the physical baryon density $\omega_{b}$ are somehow calibrated, one can easily show that an increase in $H_{0}$ must be accompanied by an increase in $\omega_{c}$. Under minimal assumptions concerning the primordial power spectrum, this inevitably leads to a potentially problematic increase in the clustering parameter $S_{8}$ Poulin et al. (2024). It so happens that $\Omega_{m}$ and $\omega_{b}$ can in fact be calibrated in a fairly robust way, through a combination of BAO and uncalibrated SNeIa in the former case, and BBN in the latter. In the rest of our work we discuss these points in more detail, explicitly demonstrating on real data how an increase in $H_{0}$ is accompanied by an increase in $\omega_{c}$ (and $S_{8}$), while studying the impact of the $\Omega_{m}$ calibration and arguing that a high-fidelity, as model-independent as possible calibration of $\Omega_{m}$, should be a major priority in the quest towards solving the Hubble tension.
@@ II Multidimensionality of the Hubble tension
$\frac{\delta\omega_{c}}{\omega_{c}}=\frac{\delta\omega_{c}}{\omega_{m}}\frac{\omega_{m}}{\omega_{c}}\approx 2\left(1+\frac{\omega_{b}+\omega_{\nu}}{\omega_{c}}\right)\frac{\delta h}{h}\,.$ (4)
@@
We remark once more that the earlier arguments hinge upon calibrations for $\Omega_{m}$ and $\omega_{b}$. We now discuss in more detail how this is achieved. To set the stage, in what follows we work within a spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric, and assume that the Etherington distance-duality relation (DDR) Etherington (1933) holds, in agreement with recent data Santos da Costa et al. (2015); Hogg et al. (2020b); Renzi et al. (2022); Liu et al. (2023); Qi et al. (2024). Within our assumptions, the transverse comoving distance $D_{M}$ is given by the following:
@@
A single BAO measurement (at a single effective redshift) of one among $\theta_{d}$, $\delta z_{d}$, or $\theta_{v}$ is unable to disentangle the effects of $r_{d}H_{0}$ and $\Omega_{m}$, which are completely degenerate between each other. However, this degeneracy can be partially broken if the BAO angular scale is measured over a sufficiently wide range of effective redshifts (at present, $0.1\lesssim z_{\text{eff}}\lesssim 2.5$). The reason is that the slope describing the $r_{d}H_{0}$-$\Omega_{m}$ correlation slowly changes with $z_{\text{eff}}$ (see e.g. Fig. 2 of Ref. Lin et al. (2021)), reflecting how the importance of the dark energy contribution relative to the matter one decreases with increasing $z_{\text{eff}}$. This is not sufficient to lead to a very constraining inference of $\Omega_{m}$ at the current level of precision (which is why earlier we stated that the degeneracy is only partially broken), but illustrates where the sensitivity of BAO measurements to $\Omega_{m}$ arises from.
@@
For what concerns the physical baryon density $\omega_{b}$, this is tightly constrained by BBN considerations on the abundance of light elements, especially deuterium. This is because $\omega_{b}$ controls (among others) the neutron-to-proton density at the time of BBN, a parameter to which the yield of light elements is particularly sensitive. It is worth noting that the CMB is also sensitive to $\omega_{b}$ through its impact on the relative height of odd and even acoustic peaks. The signature of $\omega_{b}$ in CMB data is therefore quite clean, making its determination relatively stable across different models, while still being in part model-dependent. Nevertheless, in support of the BBN determination of $\omega_{b}$, it is worth noting that this is in excellent agreement with the $\Lambda$CDM-based CMB determination of the same parameter. Moreover, it has been shown in Ref. Motloch (2020) that $\omega_{b}$ can be inferred in eight independent ways from CMB data, reflecting eight independent ways in which the physical baryon density affects the CMB power spectra, and with all eight determinations being in agreement between each other. Therefore, while the earlier calibration of $\Omega_{m}$ is somewhat model-dependent (it implicitly depends on the assumed late-time background expansion), the calibration of $\omega_{b}$ can be considered extremely robust and model-independent.
@@
where $T(k)$ denotes the transfer function, $k_{\text{eq}}$ is the equality wavenumber, and $a_{\text{eq}}$ is the equality scale factor. This necessarily increases $S_{8}\sim H_{0}^{2}$ as well, worsening the discrepancy between values of $S_{8}$ determined by weak lensing Troxel et al. (2018); Heymans et al. (2021), redshift space distortions Efstathiou and Lemos (2018); Benisty (2021); Nunes and Vagnozzi (2021), and cluster counts Sakr et al. (2022), all of which fall somewhat lower compared to the Planck (TTTEEE+lowE+lensing) determination of $S_{8}=0.832\pm 0.013$ Aghanim et al. (2020) [^3] We note that recent analyses of cosmic shear data from DES Y3 and KiDS-1000 Abbott et al. (2023), and of the cluster mass function from SRG/eROSITA Ghirardini et al. (2024), have decreased the significance of the $S_{8}$ discrepancy. Analogously to Eq. (3), we can expect the following: